# The Neglected Tails of Vision-Language Models

### 23 Jan 2024 v1

@@ 3.2 Discussions and Remarks
Human-in-the-loop validation. As VLMs’ pretraining data does not contain ground-truth concept labels, we also perform manual validation. To do so, we first construct a small validation set by downloading a balanced set of pretraining data (32 image-text pairs per concept). Then, for each concept, we tune the concept definitions for Llama-2 till reaching $>$85% retrieval precision on the validation set. In particular, since [19] releases multiple definitions per concept, we select the best ones that lead to the highest precision over the validation set. For example, the class samoyed in ImageNet refers to a dog breed; we find the definition “a breed of large white herding dog with a thick coat, native to the Ural Mountains” to be more precise than others, e.g., “a member of a people inhabiting northwestern Siberia”. To facilitate future research, we will open-source our code for LLM-based analysis and release all concept synonyms and filtered captions.
The prevalent long tails in VLMs. Our analysis reveals a notable long-tailed distribution of visual concepts (from standard benchmarks like ImageNet [9]) within widely-used pretraining datasets like LAION-400M and LAION-2B (cf. Fig. 1a). Additionally, we plot per-class zero-shot accuracies of OpenCLIP models (pretrained on LAION), establishing a strong correlation between the long-tailed distribution of concepts and the imbalanced performance of VLMs (cf. Fig. 1b). We also plot the per-class accuracies of CLIP [33] and MetaCLIP [47], which are trained on private datasets. Interestingly, they show similar imbalanced performances across ImageNet’s concepts. We observe the same long-tailed concept frequencies and imbalanced per-class accuracies across eight more benchmarks (cf. Table 7), e.g., Flowers [28] and Pets [31]. Notably, our method can find rare (tailed) concepts that challenge downstream models including the state-of-the-art GPT-4Vision [29] and DALL-E 3 [39] (cf. Fig. 8 and 9). In sum, our analysis shows the prevalent long-tail issues in VLMs.
@@
  • Internet data are naturally long-tailed: Despite balanced sampling in [47], the resulting distribution is still long-tailed. For example, while each search query is capped at 20K images, the average is around $\sim$1.2K images per query. In fact, we find that over half of the queries, such as “tailed frog” (a frog), “gyromitra” (a fungus), and “poke bonnet” (a traditional hat), have fewer than 50 samples, likely because these concepts are rare on the web.
@@ 4 Retrieval-Augmented Learning
To address the biased performance of VLMs in zero-shot recognition, we propose REtrieval-Augmented Learning (REAL), which improves performance without using any data from downstream tasks by retrieving pretraining data relevant to the target concepts. REAL has two variants: REAL-Prompt and REAL-Linear. REAL-Prompt is a novel prompting strategy that replaces the original concept name with its most frequent synonym found in the pretraining texts. REAL-Linear retrieves images relevant to the concepts from the pretraining data to form a class-balanced dataset and use this set to learn a robust linear classifier. Below we elaborate on these two methods.
@@ 4.1 REAL-Prompt for Zero-Shot Prompting
In our analysis, we discover that some synonyms for a concept might be more frequent in pretraining texts than the concept itself. Therefore, we propose using the most frequent synonym of a concept to construct prompts. Specifically, we utilize an LLM (ChatGPT [29]) to enumerate all synonyms for each concept. Next, we count their individual frequencies in the pretraining texts by string matching. We use the most frequent synonym of each concept in the prompt to construct an off-the-shelf classifier $W_{zs}$ for zero-shot recognition following CLIP [33]. This simple change leads to notably better zero-shot accuracy than using the original concept names (cf. Fig. 3) released by [33], which are hand-crafted over one year. As shown in Table 1, our approach also outperforms recent LLM-based prompting methods that use additional visual descriptors along with the given concept names (e.g., DCLIP [26] and CuPL [32]).
Synonym filtering using OpenCLIP’s text encoder. ChatGPT sometimes generate noisy synonyms. For example, for “tiger”, it lists “big cat” as a synonym, which could be easily confused with the ImageNet class “tabby cat”. To address this, we use OpenCLIP’s text encoder to filter out synonyms that might be confused with other downstream concepts. We retain only those synonyms that have the highest cosine similarity scores with their original class names. This filtering step is fully automated, ensuring a fair comparison with [26, 32] that perform LLM-based prompting without human input. We show that this step is crucial to REAL-Prompt’s performance in Table 15.
@@ 4.2 REAL-Linear for Linear Classifier Training
REAL-Linear’s exceptional efficiency. REAL-Linear is significantly more efficient than the state-of-the-art REACT [22] and can be done using only academic-scale compute resources. Unlike REACT, which requires running VLMs to extract features from all pretraining images and texts, REAL-Linear processes only pretraining texts via string matching. As a result, it can process all LAION-400M captions in one hour, whereas REACT needs 250 GPU hours just for feature extraction. In addition, REACT’s contrastive training demands extensive resources for better performance, e.g., large batch sizes (4,096) and lengthy training (up to 256 GPU hours on 16 V100 GPUs). In contrast, our linear-probing approach trains in minutes on a modest (12GB) GPU. Table 3 compares the efficiency between REACT and REAL.
@@ 5 Experimental Results
We assess REAL’s performance in zero-shot recognition tasks, comparing it against prior works on standard benchmarks. We also ablate design choices in REAL that contribute to its performance. Moreover, we show that integrating REAL with existing methods leads to improved results. Lastly, we explore an application of REAL-Prompt in text-to-image models like DALL-E 3 and SD-XL, which improves image generation of rare concepts.
@@ 5.1 Experimental setup
Datasets and metric. We report mean per-class accuracy on standard classification benchmarks, including ImageNet [9], Flowers [28], Cars [16], Aircraft [24], Pets [31], Food [5], DTD [8], EuroSAT [11], and CUB [41]. Moreover, we use variants of ImageNet to study the robustness of our methods, including ImageNet-V2 [15], ImageNet-Adversarial [13], ImageNet-Rendition [12], and ImageNet-Sketch [44]. Table 6 lists their detailed descriptions.
Compared methods. We compare against state-of-the-art zero-shot recognition methods for VLMs. We report various prompting strategies that directly use the given concept names, including prompt templates such as {concept}, “a photo of {concept}, and OpenAI’s hand-engineered templates [33]. We also compare with LLM-based prompting methods, DCLIP [26] and CuPL [32], which uses GPT to generate descriptions for constructing prompts. Finally, we compare our method with the retrieval-augmented method REACT [22]. We report two variants of REACT: Locked-Text and Gated-Image.
@@
Implementation details. In this work, we adopt a series of OpenCLIP VLMs [14, 6], which are publicly available along with their two pretraining datasets, namely LAION-400M [35] and LAION-2B [36]. We report the performance of OpenCLIP ViT-B/32 architecture in the main paper and show that REAL generalizes to other architectures in Table 13. For our method REAL-Linear that learns a linear classifier, we simply use the hyperparameters provided by prior work [20, 46]. We use a single GeForce RTX 2080 Ti (12GB) to train all the models and allocate 50GB of storage to host retrieved data for all the nine benchmark datasets.
@@ 5.2 Results
Using frequent synonyms improves zero-shot recognition. Table 1 (top half) compares REAL-Prompt (based on OpenAI’s prompt templates) against other prompting-based methods like DCLIP [26] and CuPL [32], which use GPT to generate visual descriptions. REAL-Prompt significantly outperforms them simply by replacing the original concept names with their most frequent synonyms. This underscores the importance of re-considering concept names in prompting; using the frequently used synonyms improves the recognition of related images.
@@
REAL-Linear achieves the state-of-the-art. Table 1 (bottom half) compares our REAL-Linear against the recent retrieval-augmented method REACT [22]. REAL-Linear achieves $\sim$3higher accuracy averaged across eight benchmarks than REACT, while using only 500 retrieved images per concept compared to REACT’s 10K. Importantly, REAL-Linear is significantly more efficient (cf.Table 3): it requires $\sim$1of REACT’s computes, making it more accessible to the research community.
@@
REAL benefits both head and tail classes. Table 2 shows that REAL boosts performance for both tail (least frequent 20%) and head (the rest 80%) classes on ImageNet and all nine datasets. For more specific improvements on each dataset, see Table 8.
@@
REAL benefits existing zero-shot methods. Our methods can be readily applied with existing methods to further improve performance. Table 4 shows that REAL-Prompt’s most frequent synonyms can be applied on any prompt templates, including LLM-generated ones like DCLIP [26] and CuPL [32]. Likewise, REAL-Linear can be applied on top of REACT’s finetuned OpenCLIP models, which establishes a new state-of-the-art zero-shot performance as shown in Table 5.
Improving image synthesis using REAL-Prompt. Fig. 5 shows that, while state-of-the-art generative models such as DALL-E 3 [39] and SD-XL [2] may fail to generate correct images for some rare concepts (identified by frequency estimation on LAION-400M), replacing the rare concepts used in prompts with their most frequent synonyms (found by REAL-Prompt) can help generate more accurate images. More qualitative examples can be found in Fig. 10 and 11.
@@
Ablation studies for REAL-Linear. Our experiments shed light on key design aspects of REAL-Linear, with major insights summarized below. Table 9 shows that using all concept synonyms, as opposed to just the original concept names, can help retrieve more diverse pretraining data, improving the averaged accuracy by 4%. Table 11 demonstrates that learning the linear classifier with both text and image features using cross-modal WiSE-FT [20, 46] leads to a 6.4% increase compared to linear probing with only image features. Lastly, Table 12 shows that increasing the retrieval size from 100 to 500 per concept yields a modest accuracy improvement of 0.9%. Based on this, we adopt 500 as the standard retrieval size for our experiments.
@@
Table 3: Compute cost comparison between REACT [22] and our REAL-Linear. We ablate resources required for each method on the ImageNet experiment. Clearly, REAL-Linear (retrieving 500 pretraining images per concept) uses much less compute than REACT, e.g., retrieving 1000$\times$ less images, using 400$\times$ less storage and 10,000$\times$ less training time.
@@
GPU memory 256 GB 3 GB 1%
Figure 5: Improving image generation using REAL-Prompt. We show two rare concepts identified by our frequency estimation where DALL-E 3 and SD-XL struggle to generate accurate images: bank swallow (top, a bird from the CUB dataset) and thorn apple (bottom, a flower from the Flowers dataset). Using their original names, both DALLE 3 and SD-XL incorrectly render the bird’s colors. Additionally, DALLE 3 erroneously adds thorns to the flower, while SD-XL depicts an apple with literal thorns. Instead, using the most frequent synonyms (sand martin for bank swallow, datura for thorn apple) as found by REAL-Prompt results in both systems generating accurate images. See more examples in Fig. 10 and 11.
Table 4: Improvements using REAL-Prompt with existing prompting methods. Our REAL-Prompt can be combined with existing prompt templates including OpenAI’s hand-engineered templates and LLM-generated templates [33] like DCLIP [26] and CuPL [32]. On datasets such as Flowers, DTD, and EuroSAT, integrating REAL-Prompt results in an accuracy boost of 5$\sim$8%.
@@ 5.3 Discussions
Broad impacts. Our work has positive societal impacts. Our concept frequency estimation method not only confirms the long-tailed distribution of concepts in VLMs’ pretraining data but also reveals a strong correlation with VLMs’ imbalanced or biased performances. By identifying concepts that VLMs insufficiently learn, we can implement targeted measures to prevent unfair or biased predictions related to these concepts.
Limitations and future work. We acknowledge several limitations in our methods. First, while we offer a method to estimate concept frequency, we cannot accurately evaluate its precision and recall due to the absence of ground-truth annotations of the pretraining data. Second, this approach relies only on textual captions and may overlook the broader concept coverage that images offer. Lastly, filtering ambiguous captions using off-the-shelf LLMs for each caption-concept pair is time-consuming. We expect the future work to address these limitations.
@@ 6 Conclusions
We explore the neglected long-tail issues existing in Vision-Language Models (VLMs). We present a novel method for estimating concept frequency in VLMs’ large-scale multimodal pretraining data. With our method, we reveal that concepts defined by downstream tasks (e.g., image classification) follow long-tailed distributions in the pretraining data. Importantly, for the first time, we establish the strong correlation between the long-tailed concept distributions and the imbalanced or biased performance (e.g., per-class accuracy in zero-shot recognition) of VLMs. Further, to mitigate VLMs’ imbalanced performance through the lens of zero-shot recognition, we propose retrieval-augmented learning (REAL), consisting of two variants: REAL-Prompt and REAL-Linear. First, instead of using the original class names defined by a downstream task, REAL-Prompt adopts the corresponding most frequent synonyms found in the pretraining texts. This simple change already outperforms human-engineered and LLM-generated prompts across nine standard benchmarks. Second, REAL-Linear uses all the synonyms to retrieve a small class-balanced set of data from the pretraining dataset and trains a robust linear classifier atop the VLM. This approach outperforms the recent retrieval-augmented solution REACT, using 400×400$\times$400 × less storage and 10,000$\times$ less training time. Finally, we show that contemporary text-to-image generators (e.g., DALL-E 3 and SD-XL) may fail to generate images featuring rare concepts. As such, we apply REAL-Prompt to these generators and show that using most frequent synonyms of a rare concept can help them synthesize more accurate images.

### 02 Feb 2024 v2

@@ 3.2 Discussions and Remarks
Human-in-the-loop validation. As VLMs’ pretraining data does not contain ground-truth concept labels, we perform manual validation to ensure LLM performs well in filtering the irrelevant texts. To do so, we first construct a small validation set by downloading a balanced set of pretraining data (32 image-text pairs per concept). Then, for each concept, we tune the concept definitions for Llama-2 till reaching $>$85% retrieval precision on the validation set. In particular, since [21] releases multiple definitions per concept, we select the best ones that lead to the highest precision over the validation set. For example, the class samoyed in ImageNet refers to a dog breed; we find the definition “a breed of large white herding dog with a thick coat, native to the Ural Mountains” to be more precise than others, e.g., “a member of a people inhabiting northwestern Siberia”. To facilitate future research, we will open-source our code for LLM-based analysis and release all concept synonyms and filtered captions.
The prevalent long tails in VLMs. Our analysis reveals an ever-neglected long-tailed distribution of visual concepts (from standard benchmarks like ImageNet [10]) within widely-used pretraining datasets like LAION-400M and LAION-2B (cf. Fig. 1a). Additionally, we plot per-class zero-shot accuracies of OpenCLIP models (pretrained on LAION), establishing a strong correlation between the long-tailed distribution of concepts and the imbalanced performance of VLMs (cf. Fig. 1b). We also plot the per-class accuracies of CLIP [35] and MetaCLIP [49], which are trained on private datasets. Interestingly, they show similar imbalanced performance across ImageNet’s concepts, likely because Internet data follows a similar long-tailed distribution. We observe the same trend across eight more benchmarks (cf. Table 7), e.g., Flowers [30] and Pets [33]. Notably, our frequency estimation method helps find rare (tailed) concepts that challenge popular multimodal systems including the state-of-the-art GPT-4Vision [31] and DALL-E 3 [41] (cf. Fig. 8 and 9). In sum, our analysis shows that long-tailed issues are prevalent in VLMs.
@@
  • Internet data are naturally long-tailed: Despite balanced sampling in [49], the resulting distribution is still long-tailed. For example, while each search query is capped at 20K images, the average is around $\sim$1.2K images per query. In fact, we find that over half of the queries, such as “tailed frog” (a frog), “gyromitra” (a fungus), and “poke bonnet” (a traditional hat), have less than 50 samples, likely because these concepts are rare on the web.
@@ 4 Retrieval-Augmented Learning
To address the biased performance of VLMs in zero-shot recognition, we propose REtrieval-Augmented Learning (REAL), which improves performance without using any data from downstream tasks by retrieving pretraining data relevant to the target concepts. REAL has two variants: REAL-Prompt and REAL-Linear. REAL-Prompt is a novel prompting strategy that replaces the original concept name with its most frequent synonym found in pretraining texts. REAL-Linear retrieves images relevant to the concepts from the pretraining data to form a more balanced subset for training a robust linear classifier. Below we elaborate on these two methods.
@@ 4.1 REAL-Prompt for Zero-Shot Prompting
In our analysis, we discover that some synonyms for a concept might appear more frequently in pretraining texts than the concept itself. Therefore, we propose using the most frequent synonym of a concept to construct prompts. Specifically, we utilize an LLM (ChatGPT [31]) to enumerate all synonyms for each concept. Next, we count their individual frequencies in the pretraining texts by string matching. We use the most frequent synonym of each concept in the prompt to construct an off-the-shelf classifier $W_{zs}$ for zero-shot recognition following CLIP [35]. This simple change leads to significantly better zero-shot accuracy than using the original concept names (cf. Fig. 3) released by [35], which are hand-crafted over one year. As shown in Table 1, our approach also outperforms recent LLM-based prompting methods that use additional visual descriptors along with the given concept names (e.g., DCLIP [28] and CuPL [34]).
Synonym filtering using OpenCLIP’s text encoder. ChatGPT sometimes generates noisy synonyms. For example, for “tiger”, it lists “big cat” as a synonym, which could be easily confused with another ImageNet class “tabby cat”. To address this, we use OpenCLIP’s text encoder to filter out synonyms that might be confused with other downstream concepts. We retain only those synonyms that have the highest cosine similarity scores with their original class names. This filtering step is fully automated, ensuring a fair comparison with [28, 34] that perform LLM-based prompting without human input. We show that this step is crucial to REAL-Prompt’s performance in Table 15.
@@ 4.2 REAL-Linear for Linear Classifier Training
REAL-Linear’s exceptional efficiency. REAL-Linear is significantly more efficient than the state-of-the-art REACT [24] and can be done using only academic-scale computing resources. Unlike REACT, which requires downloading the whole pretraining dataset and running VLMs to extract features for all pretraining images and texts, REAL-Linear processes only pretraining texts via string matching. As a result, it can process all LAION-400M captions in one hour, whereas REACT needs 250 GPU hours just for feature extraction. In addition, REACT’s contrastive training demands extensive resources for better performance, e.g., large batch sizes (4,096) and lengthy training (up to 256 GPU hours on 16 V100 GPUs). In contrast, our linear-probing approach trains in minutes on a modest (12GB) GPU. Table 3 compares the efficiency between REACT and REAL.
@@ 5 Experimental Results
We show that REAL establishes the new state-of-the-art for zero-shot recognition, outperforming existing prompting and retrieval-augmented methods across standard benchmarks. We ablate the design choices of REAL, revealing technical insights that contribute to its superior performance. We show that REAL can be combined with existing methods for even better results. Moreover, we show that REAL-Prompt improves image generation of rare concepts using text-to-image models like DALL-E 3 and SD-XL.
@@ 5.1 Experimental setup
Datasets and metric. We report mean per-class accuracy on standard classification benchmarks, including ImageNet [10], Flowers [30], Cars [18], Aircraft [26], Pets [33], Food [5], DTD [9], EuroSAT [13], and CUB [43]. Moreover, we use variants of ImageNet to study the out-of-distribution (OOD) robustness of our methods, including ImageNet-V2 [17], ImageNet-Adversarial [15], ImageNet-Rendition [14], and ImageNet-Sketch [46]. Table 6 details these datasets.
Compared methods. We compare against state-of-the-art zero-shot recognition methods for VLMs. We report various prompting strategies that directly use the given concept names, including prompt templates such as {concept}, “a photo of {concept}, and OpenAI’s hand-engineered templates [35]. We also compare with LLM-based prompting methods, DCLIP [28] and CuPL [34], which uses GPT to generate visual descriptions for constructing prompts. Finally, we compare our method with the retrieval-augmented SOTA REACT [24] that retrieves data using VLMs’ features and performs contrastive finetuning. We report two variants of REACT: Locked-Text and Gated-Image.
@@
Implementation details. In this work, we ablate on a series of OpenCLIP VLMs [16, 7], which are publicly available along with their two pretraining datasets, namely LAION-400M [37] and LAION-2B [38]. We report the performance of OpenCLIP ViT-B/32 architecture in the main paper and show that REAL generalizes to other architectures in Table 13. For our REAL-Linear that learns a linear classifier, we simply use the hyperparameters provided by prior work [22, 48]. We use a single GeForce RTX 2080 Ti (12GB) to train all the models and allocate 50GB of storage to host retrieved data for all the nine benchmark datasets.
@@ 5.2 Results
Using frequent synonyms improves zero-shot recognition. Table 1 (top half) compares REAL-Prompt (based on OpenAI’s prompt templates) against other prompting-based methods like DCLIP [28] and CuPL [34], which use GPT to generate visual descriptions. REAL-Prompt significantly outperforms them simply by replacing the original concept names with their most frequent synonyms. This highlights the need to reconsider concept names in prompts. REAL-Prompt is also much cheaper because it queries ChatGPT for synonyms, which are shorter than the rich visual descriptions queried by DCLIP and CuPL. For every 1K concepts, DCLIP and CuPL generate 50K (\$currency-dollar\\$\$0.5) and 500K (\$currency-dollar\\$\$5) tokens using ChatGPT, respectively, while REAL-Prompt only requires 10K tokens (\$currency-dollar\\$\$0.1).
@@
REAL-Linear achieves the state-of-the-art. Table 1 (bottom half) compares our REAL-Linear against the retrieval-augmented SOTA REACT [24]. REAL-Linear achieves $\sim$3higher accuracy averaged across eight benchmarks than REACT, while using only 500 retrieved images per concept compared to REACT’s 10K. Importantly, REAL-Linear is significantly more efficient (cf. Table 3): it requires $\sim$1of REACT’s computes, making it more accessible to the research community. It also outperforms another recent method NeuralPriming [44] under its experimental setup (cf. Table 10).
@@
REAL improves accuracy on tail classes. Table 2 shows that REAL boosts performance for both tail (least frequent 20%) and head (the rest 80%) classes on ImageNet and all nine datasets. For more specific improvements on each dataset, see Table 8.
@@
REAL benefits existing zero-shot methods. Our methods can be readily applied with existing methods to further improve performance. Table 4 shows that REAL-Prompt’s most frequent synonyms can be applied on any prompt templates, including LLM-enriched ones like DCLIP [28] and CuPL [34]. Likewise, REAL-Linear can be applied on top of REACT’s finetuned OpenCLIP models, which achieves even better performance as shown in Table 5.
@@
Ablation studies for REAL-Linear. To understand REAL-Linear’s superior performance, we conduct several experiments, with key insights summarized below. Table 9 shows that using all concept synonyms, as opposed to just the original concept names, can help retrieve more diverse pretraining data, improving the averaged accuracy by 4%. Table 11 demonstrates that learning the linear classifier with both text and image features using cross-modal WiSE-FT [22, 48] leads to a 6.4% increase compared to linear probing with only image features. Lastly, Table 12 shows that increasing the retrieval size from 100 to 500 per concept yields a modest accuracy improvement of 0.9%. Based on this, we adopt 500 as the standard retrieval size for our experiments.
Improving image synthesis using REAL-Prompt. Fig. 5 shows that, while state-of-the-art generative models such as DALL-E 3 [41] and SD-XL [2] may fail to generate correct images for some rare concepts (identified by our frequency estimation on LAION-400M), replacing the rare concepts used in prompts with their most frequent synonyms (found by REAL-Prompt) can help generate more accurate images. More qualitative examples can be found in Fig. 10 and 11.
@@
Table 3: Compute cost comparison between REACT [24] and our REAL-Linear. We compare the resources required for each method on the ImageNet experiment. Clearly, REAL-Linear (retrieving 500 pretraining images per concept) uses much less compute than REACT, e.g., retrieving 1000$\times$ less images, using 400$\times$ less storage and 10,000$\times$ less training time.
@@
GPU memory 256 GB 2 GB 0.8%
Figure 5: Improving image generation using REAL-Prompt. We show two rare concepts identified by our frequency estimation where DALL-E 3 and SD-XL struggle to generate accurate images: bank swallow (top, a bird from the CUB dataset) and thorn apple (bottom, a flower from the Flowers dataset). Using their original names, both DALL-E 3 and SD-XL incorrectly render the bird’s colors. Additionally, DALL-E 3 erroneously adds thorns to the flower, while SD-XL depicts an apple with literal thorns. Instead, using the most frequent synonyms (sand martin for bank swallow, datura for thorn apple) as found by REAL-Prompt results in both systems generating accurate images. See more examples in Fig. 10 and 11.
Table 4: Improvements using REAL-Prompt with existing prompting methods. Our REAL-Prompt can be combined with existing prompt templates including OpenAI’s hand-engineered templates and LLM-enriched templates [35] like DCLIP [28] and CuPL [34]. On datasets such as Flowers, DTD, and EuroSAT, integrating REAL-Prompt results in an accuracy boost of 5$\sim$8%.
@@ 5.3 Discussions
Broad impacts. Our work has positive societal impacts. Our concept frequency estimation method explains why VLMs are biased (or imbalanced) by confirming the long-tailed distribution of concepts in their pretraining data. By identifying concepts that VLMs have insufficiently learned, we can implement targeted measures to prevent unfair or biased predictions related to these concepts.
Limitations and future work. We acknowledge several limitations in our methods. First, while we offer a method to estimate concept frequency, we cannot accurately evaluate its precision and recall due to the absence of ground-truth annotations of the pretraining data. Second, our estimation method relies only on the textual captions which may overlook other visual concepts that are present in the images but not in their captions. Lastly, filtering ambiguous captions using off-the-shelf LLMs for each caption-concept pair is time-consuming. We expect the future work to address these limitations.
@@ 6 Conclusions
We investigate the critical yet ever-neglected long-tailed issues of Vision-Language Models (VLMs). We use large language models (LLMs) to estimate concept frequency in VLMs’ large-scale multimodal pretraining data, uncovering long-tailed distributions for concepts used in downstream tasks. Crucially, we demonstrate a strong correlation between the long-tailed concept distributions and VLMs’ imbalanced zero-shot performance. To address this imbalance, we propose retrieval-augmented learning (REAL), with two variants: REAL-Prompt and REAL-Linear. REAL-Prompt replaces original class names from downstream tasks with their most common synonyms found in the pretraining texts, outperforming both human-engineered and LLM-enriched prompts. On the other hand, REAL-Linear leverages concepts synonyms to fetch a balanced subset of pretraining data for training a robust linear classifier atop the frozen VLM, surpassing the previous SOTA using 400×400$\times$400 × less storage and 10,000$\times$ less training time. Finally, we highlight that modern text-to-image generators (e.g., DALL-E 3 and SD-XL) often fail to generate images for the rare concepts identified by our frequency estimation method. By applying REAL-Prompt, we demonstrate that using common synonyms helps generate more accurate images.